# \$*\$ \$*\$ footnotetext: Corresponding Author

### v1

@@ 2.3 Unified Multimodal Models
An alternative and promising direction [88, 68, 33, 37, 91, 74, 16] for unifying multimodal understanding and generation involves assembling off-the-shelf specialized LMMs and visual generative models by tuning adapters or learnable tokens. Representative works [103, 37, 74, 16] have demonstrated the promissing capabilities and intriguing properties of such assembled unified frameworks, highlighting their potential for further exploration.
@@ 4.1 Experimental Setup
Datasets. The curated approximately 66M image-text pairs consist of images with a resolution of at least 512 pixels in width and height. The images are filtered from CC12M [12], COYO [11], LAION-Aesthetic-12M∗*∗ ∗*∗ * https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap and AI synthetic data. The images are recaptioned by ShareGPT4-V [18] except for the synthetic data. The 9M high-quality multimodal understanding instruction data is curated from Densefusion-1M [57], and LLaVA-OneVision [51].
@@
Table 3: Evaluation on multimodal understanding benchmarks. # Params. indicates the number of parameters of base LLM. * indicates the method uses two distinct models or sets of parameters for multimodal understanding and generation, respectively. Und. indicates “understanding”.
@@
Native Unified Show-o [108] 1.3B 1097.2 58.0 51.5 - 27.4 - -
@@ 4.2 Multimodal Understanding
Table 4: Evaluation on the GenEval [38] benchmark. Gen. denotes “generation”. # Params. indicates the number of parameters of base LLM. # Data. indicates the number of image-text pairs used for visual generation during training. * indicates the method uses two distinct models or sets of parameters for multimodal understanding and generation, respectively. Obj.: Object. Attri.: Attribute. Our results are obtained based on the rewritten dense prompts.
@@
Gen. Only SD3 (d=24) [34] - 0.98 0.74 0.63 0.67 0.34 0.36 0.62
SD3-Medium [34] - - 0.99 0.94 0.72 0.89 0.33 0.60 0.74
Unifying via SEED-X [37] 17B 158M 0.97 0.58 0.26 0.80 0.19 0.14 0.49
Tailored TokenFlow-XL [79] 14B 60M 0.95 0.60 0.41 0.81 0.16 0.24 0.55
Models ILLUME [96] 7B 15M 0.99 0.86 0.45 0.71 0.39 0.28 0.61
@@ Appendix A Technical Appendices and Supplementary Material
Table 11: Comparative analysis of selected unified multimodal models based on the utilization of visual representations and type of unified modeling for multimodal understanding and generation. In this context, native und. & gen. refers to the direct decoding of output sequences into texts and images, as opposed to serving as conditions for decoding using external pre-trained decoders like Stable Diffusion. Please refer to the complete table in the appendix. indicates the method uses two distinct models for multimodal understanding and generation, respectively.

### v2

@@ 2.3 Unified Multimodal Models
An alternative and promising direction [94, 71, 34, 38, 98, 77, 17] for unifying multimodal understanding and generation involves assembling off-the-shelf specialized LMMs and visual generative models by tuning adapters or learnable tokens. Representative works [111, 38, 77, 17] have demonstrated the promising capabilities and intriguing properties of such assembled unified frameworks, highlighting their potential for further exploration.
@@ 4.1 Experimental Setup
Datasets. The curated approximately 66M image-text pairs consist of images with a resolution of at least 512 pixels in width and height. The images are filtered from CC12M [13], COYO [12], LAION-Aesthetic-12M∗*∗ ∗*∗ * https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap and AI synthetic data. The images are recaptioned by LMMs except for the synthetic data. The 9M high-quality multimodal understanding instruction data is curated from Densefusion-1M [58], and LLaVA-OneVision [52].
@@
Table 3: Evaluation on multimodal understanding benchmarks. # Params. indicates the number of parameters of base LLM. * indicates the method uses two distinct models or sets of parameters for multimodal understanding and generation, respectively. Und. indicates “understanding”. Results in gray indicate the performance of und. only models or models with total parameters more than 13B.
@@
Native Unified BAGEL [30] 14B 1687.0 - - 85.0 55.3 - -
Show-o [118] 1.3B 1097.2 58.0 51.5 - 27.4 - -
@@ 4.2 Multimodal Understanding
Table 4: Evaluation on the GenEval [39] benchmark. Gen. denotes “generation”. # Params. indicates the number of parameters of base LLM. # Data. indicates the number of image-text pairs used for visual generation during training. * means the method uses two distinct models for multimodal understanding and generation, respectively. Obj.: Object. Attri.: Attribute. Our results are obtained using rewritten prompts. + indicates the additional data required by the pretrained diffusion models.
@@
Gen. Only SD3-Medium [35] - - 0.99 0.94 0.72 0.89 0.33 0.60 0.74
Unifying via SEED-X [38] 17B 158M+ 0.97 0.58 0.26 0.80 0.19 0.14 0.49
Assembling TokenFlow-XL [82] 14B 60M 0.95 0.60 0.41 0.81 0.16 0.24 0.55
Tailored ILLUME [103] 7B 15M+ 0.99 0.86 0.45 0.71 0.39 0.28 0.61
Models MetaQuery-XL [77] 7B 28M+ - - - - - - 0.80
@@ Appendix A Technical Appendices and Supplementary Material
Table 11: Comparative analysis of selected unified multimodal models based on the utilization of visual representations and type of unified modeling for multimodal understanding and generation. In this context, native und. & gen. refers to the direct decoding of output sequences into texts, images, and videos, as opposed to serving as conditions for decoding using external pre-trained decoders like Stable Diffusion. indicates the method uses two distinct models for multimodal understanding and generation, respectively.